You have 5 ul of a pure (undiluted) culture at a concentration of 3.6 x 106 cells/ml. You then add 495 ul sterile diluent. What is the dilution factor ? What is the final concentration (following the dilution) ? NOTE: Use the formula V1 x D1 = V2 x D2 to solve this problem.
Answer #1

DF = 495 + 5 / 5 = 500/5 = 100

Dilution = 1/100 = 0.01

Final concentration = 3.6 × 10^6 / 100 = 3.6 × 10^4 CFU per ml

Or

V1D1 = V2D2

(3.6 × 10^6)(5) = D2(500)

D2 = 3.6 × 10^4 CFU per ml

Please rate.